OpenStudy (anonymous):

4x^2 - 16 = 0 2x^2 = 16 x^2 + 5 = 0 can someone plzz solve these equations for me plzzz

OpenStudy (amistre64):

we are not here to do the work for you, but rather to guide you along the process

OpenStudy (anonymous):

okay so help me

OpenStudy (amistre64):

these are pretty basic, just one term containing the variable; algebra everything else to the other side of the equal sign; and sqrt

OpenStudy (anonymous):

im still confused

OpenStudy (amistre64):

how do you work an equation?

OpenStudy (amistre64):

4u - 16 = 0, solve for u

OpenStudy (anonymous):

subtract or add by both side idk

OpenStudy (amistre64):

thats a good start yes :) so lets add 16 to each side, then divide off that 4

OpenStudy (anonymous):

okay let me try :)

OpenStudy (anonymous):

so 4u=10 and then divide

OpenStudy (amistre64):

4u - 16 = 0 +16 +16 ------------ 4u = 16 but yes, then divide by 4

OpenStudy (anonymous):

lol sorry how did i get 10 i ment 16 sorry

OpenStudy (anonymous):

so the answer is 4

OpenStudy (amistre64):

the answer is almost ready :) one more thing to consider

OpenStudy (anonymous):

so wat about 2x^2=16

OpenStudy (anonymous):

yeah wat ?

OpenStudy (amistre64):

instead of naming the unknown as "u", lets name it as "x^2" x^2 = 4 ; now we to undo that ^2 part

OpenStudy (amistre64):

take the square root of each side \[\sqrt{x^2}=\pm\sqrt{4}\] \[x=\pm2\]

OpenStudy (anonymous):

okay i get that part

OpenStudy (amistre64):

2x^2 = 16, dont let that x^2 scare you, just rename it if its bothering you: say x^2 = u 2u = 16 and solve for u

OpenStudy (anonymous):

okay ley me try

OpenStudy (anonymous):

i got 8

OpenStudy (amistre64):

good, then: u = 8 or rewritten in x^2 stuff: x^2 = 8, sqrt each side

OpenStudy (anonymous):

is it \[\sqrt{x ^{2}}and \sqrt{2}\]

OpenStudy (amistre64):

i see that your trying to look ahead perhaps :) \[x^2 = 8\] \[\sqrt{x^2} = \pm\sqrt{8}\] \[x = \pm\sqrt{4*2}\] \[x = \pm\sqrt{4}*\sqrt{2}\] \[x = \pm2\sqrt{2}\]

OpenStudy (amistre64):

the next one: x^2 + 5 = 0 works the same way, but the solution depends on what you are looking for.

OpenStudy (anonymous):

lol oh okay let me try

OpenStudy (anonymous):

so i can put it like 2u+5=o

OpenStudy (amistre64):

x^2 + 5 = 0 ; let u = x^2 and rewrite u + 5 = 0

OpenStudy (anonymous):

so u is by it self or with the 5

OpenStudy (amistre64):

theres a nice little trick i just saw, but it might confuse you a little: if we let x^2 = 5u we get: 5u + 5 = 0, solves to u=-1, x^2 = 5(-1) = -5 is just as much a valid process as the rest of it

OpenStudy (anonymous):

lol yeah it did kinda confused me

OpenStudy (anonymous):

so do i solve it like that

OpenStudy (amistre64):

lol, then lets just keep it as:\[x^2+5=0~:~let~x^2=u\] \[u+5 = 0\\~~-5~~-5\\-----\\~u~~~~~ = -5\] \[since~u=x^2, ~rewrite~as \] \[x^2=-5\]\[\sqrt{x^2}=\pm\sqrt{-5}\]\[x=\pm\sqrt{-5}\]

OpenStudy (amistre64):

have you come across anthing called imaginary numbers, or complex numbers?

OpenStudy (anonymous):

um i dont think so

OpenStudy (amistre64):

if you havent, then the answer to: x^2 + 5 = 0 is, no real solutions

OpenStudy (amistre64):

the square root of a negative number has no real solutions; it only has complex (or imaginary) solutions

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i really need to practice more math is my weakness

OpenStudy (amistre64):

practice does help :) good luck

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

can there be a square root for -25?